OpenStudy (anonymous):

Given f(x)=2x-3 and g(x)=4, for what values of x is f(x)+g(x) = f(x) ∙ g(x)?

OpenStudy (bahrom7893):

f(x)+g(x) = (2x-3) + (4)

OpenStudy (bahrom7893):

is it f(x) times g(x) or is it a circular hollow dot, like small o ?

OpenStudy (anonymous):

yes f(x) times g(x)

OpenStudy (bahrom7893):

f(x)*(gx_ = (2x-3)*4

OpenStudy (bahrom7893):

f(x)+g(x) = f(x)*g(x) (2x-3) + (4) = (2x-3)*4

OpenStudy (bahrom7893):

2x +1 = 8x - 12

OpenStudy (bahrom7893):

2x - 8x = -12 - 1 6x = -13 x = -13/6 just check my arithmetic.

OpenStudy (anonymous):

i think im supposed to find what number for x will make both sides equal?

OpenStudy (bahrom7893):

yea well we just found it, x = -13/6

OpenStudy (anonymous):

hmm ok thanks!